Grade 11Physical Chemistry

To an evacuated vessel with movable piston under external pressure of 1 atm, 0.1 mol of He and 1 mol of an unknown compound(vapour pressure=0.68 atm at 0 C) are introduced.Considering the ideal gas behaviour the total volume of the gases al 0C is

ashu
12 Years agoGrade 11

1 Answer

Amit Jain
12 Years ago
P(unknown)= P-vapour pressure=1-.68=.32
Now v= v1+v2
applying V=nRT/P
here for v1(for He) P=1 atm
and v2(for unknown gas)= .32
and solving V=n1* RT/P1+ n2* RT/ P2 we get answer

Thanks & Regards
Name,
askIITians Faculty
Qualification.
College graduated